Insert a random 9-mer containing at least one of each nucleobase in the following sequence: GCCUGUUGG-(N9)-CCAACAAU

(a) Calculate the molecular weight, percent GC content and molar absorptivity for the oligomer you designed.

(b) Provide a structural representation of your oligo using one letter nucleotide abbreviations. Show any intramolecular interactions you anticipate.

(c) Assume that fluorescein and tetramethylrhodamine were successfully bioconjugated to the 5’ and 3’ ends, respectively. Explain how the addition of a ssDNA oligo with a complementary base sequence to your 9-mer would confirm or refute your proposed structure. Use labelled illustrations to demonstrate this interaction i.e. show the hybrid and provide the sequence of the DNA strand.

(d) Suppose the hybrid was briefly subjected to N-cyclohexyl-3-aminopropanesulfonic acid (CAPS). What would this yield? Explain and demonstrate.
Answer #1

a) In this oligos , i am including 9 mer base sequence

AUCCGUUAC, so total sequence

GCCUGUUGG AUCCGUUAC CCAACAAU

Molecular weight of single stranded RNA sequence is given by

(An x 329.2) +(Un x 306.2) + ( Cn x 305.2) + ( Gn x 345.2) + 159

An, Gn , Cn and Un are number of each respective nucleotide within the oligomer

159 is a molecular weight of 5' triphosphate

Here = An = 6, Un = 7, Gn = 5 and Cn = 8

So, molecular weight is 8445 g/ mol

Gc content of oligomer determine in two steps

First tally the number of guanine or cytosine nucleotide in sequence

Then divide number of cytosine and guanine nucleotide by total number of base pair in the sequence

G+C = 5+8= 13

Total nucleotide = 26

13/26 = 1/2* 100 = 50%

% GC content = 50%

Molecular absorptivity is a molar extinction cofficient , is a measure of how well a chemical species absorb a given wavelength of light

Molar absorptivity = 234800 L/mol cm

b) 5' GCCUGUUGGAUCCGUUACCCAACAAU-3'

Their is formation of self dimer and hair pin loop

6)isi hin o Ca レ 3

c) here, the oligonucleotide is 5' and 3' modified with dye, where fluorescein and tetranethylrhodamin act as donar and acceptor dye. It is used as fluorescence resonance energy transfer. When oligonucleotide hybridize with ssDNA , if complementry sequence not avavilable then fluorecsence of the donar is visible. If target is present, labelled probe hybridize with target and FRET can occur in which donar trasfer energy to acceptor without emission of photon. Here, also the presence of fluorescence confirm or refute our proposed structure.